OpenStudy (anonymous):

Sand is being emptied from conveyor belt onto the ground. The sand is piling up in the shape of a cone. Sand is currently being added to the pile at a rate of 50ft3 / sec. At what rate is the height of the pile increasing when there is 1200ft 3 of sand, and the radius is 10 feet and increasing at 0.15 ft/sec?

OpenStudy (anonymous):

1200 cubic feet = volume of sand radius of cone = 10 rate of increase of radius = \(\Large{dr\over dt}\)=0.15 rate of volume increase = \(\Large {dV\over dt}\)=50

OpenStudy (anonymous):

\[V={1\over 3}\pi r^2h\\ {dV\over dt}=\left({1\over3}\pi\times2r{dr\over dt}\right)h+\left({1\over3}\pi r^2\right){dh\over dt} \]

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

did you follow the derivative?

OpenStudy (anonymous):

\[(\frac{ 1 }{ 3}\pi \times2r0.15)h+(\frac{ 1 }{ 3}\pi r^2)\frac{ dh }{ dt }\]yeah I got

OpenStudy (anonymous):

great.. we are asked to find \(\Large{dh\over dt}\).. plug in the numbers that I extracted from the given information

OpenStudy (anonymous):

you are missing the \[dV\over dt\] on the left side.. we need it

OpenStudy (anonymous):

50= then what I wrote

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

ok I got that part what do I do next

OpenStudy (anonymous):

plug in the numbers and solve for \[dh\over dt\]

OpenStudy (anonymous):

ok but what does r=

OpenStudy (anonymous):

r = radius of the cone

OpenStudy (anonymous):

all the values are up there except for one

OpenStudy (anonymous):

ok what does h=

OpenStudy (anonymous):

exactly \[V={1\over3}\pi r^2h\\ 1200={1\over3}\pi(10)^2h\]solve for h

OpenStudy (anonymous):

ok my calculator is not letting me calculate can yo help me calculate\[\frac{ 1 }{ 3 }\pi x 2 (10)(0.15)11.46\frac{ 1 }{ 3 }\pi (10)^2\]

OpenStudy (anonymous):

try the wolframalpha.com

OpenStudy (anonymous):

wait I thi nk I got it

OpenStudy (anonymous):

but check with the +, - signs

OpenStudy (anonymous):

great

OpenStudy (anonymous):

ok I got\[\frac{ dh }{ dt }=.01326\] is that right

OpenStudy (anonymous):

well, I did not calculate myself but thats how you gotta do it :)

OpenStudy (anonymous):

is that it or is there more

OpenStudy (anonymous):

well, that is all that the question is askin, aint it

OpenStudy (anonymous):

yess I was just making sure

OpenStudy (anonymous):

lets not give more than what they ask for

OpenStudy (anonymous):

thanks for all your help

OpenStudy (anonymous):

yer welcome.